Please show TI 83 steps to figure this out. I've written wrong answers so you can see the full questions.

In a study of government financial aid for college students, it becomes necessary to estimate the percentage of full-time college students who earn a bachelors degree in four years or less. Find the sample size needed to estimate that percentage. Use a 0.04 margin of error and use a confidence level of 90%. Complete parts (a) through (c) below. a. Assume that nothing is known about the percentage to be estimated. n 423 (Round up to the nearest integer.) b. Assume prior studies have shown that about 55% of ful-time students earn bachelors degrees in four years or less n 419 Round up to the nearest integer.) c. Does the added knowledge in part (b) have much of an effect on the sample size? O A. No, using the additional survey information from part (b) only slightly reduces the sample size B. ° C. O D. No, using the additional survey information from part (b) does not change the sample size. Yes, using the additional survey information from part (b) dramatically reduces the sample size. Yes, using the additional survey information from part (b) only slightly increases the sample size.
Answer #1

Ans:

critical z value=1.64

a)

Use invNorm function of Ti-83 calculator.

Sample size required,n=1.64^2*0.5*(1-0.5)/0.04^2=420.25

n\approx 420

b)

Sample size required,n=1.64^2*0.55*(1-0.55)/0.04^2=416.05

n\approx 416

c)Sample size slightly reduces the sample size.

Option A is correct.